find the equation of the line that is parallel to the graph of 3x-5y=3 and passes through (-6,-6). write the equation in standard form
Found 2 solutions by josgarithmetic, lwsshak3:
Answer by josgarithmetic(39620)
Answer by lwsshak3(11628)

find the equation of the line that is parallel to the graph of 3x-5y=3 and passes through (-6,-6). write the equation in standard form
***
standard form of equaation for a line: y=mx+b, m=slope, b=y-intercept
given equation:
3x-5y=3
5y=3x-3
y=3x/5-3/5
slope=3/5
..
parallel lines have the same slope
equation of parallel line
y=3x/5+b
solve for b using coordinates of given point on the line(-6,-6)
-6=3*(-6)/5+b
b=-6+18/5=-12/5
equation: y=3x/5-12/5
5y=3x-12
in standard form:
3x-5y=12